### v1

# A Novel Benchmark for Few-Shot Semantic Segmentation in the Era of Foundation Models

Abstract
In recent years, the rapid evolution of computer vision has seen the emergence of various vision foundation models, each tailored to specific data types and tasks. While large language models often share a common pretext task, the diversity in vision foundation models arises from their varying training objectives. In this study, we delve into the quest for identifying the most effective vision foundation models for few-shot semantic segmentation, a critical task in computer vision. Specifically, we conduct a comprehensive comparative analysis of four prominent foundation models: DINO V2, Segment Anything, CLIP, Masked AutoEncoders, and a straightforward ResNet50 pre-trained on the COCO dataset. Our investigation focuses on their adaptability to new semantic segmentation tasks, leveraging only a limited number of segmented images. Our experimental findings reveal that DINO V2 consistently outperforms the other considered foundation models across a diverse range of datasets and adaptation methods. This outcome underscores DINO V2’s superior capability to adapt to semantic segmentation tasks compared to its counterparts. Furthermore, our observations indicate that various adapter methods exhibit similar performance, emphasizing the paramount importance of selecting a robust feature extractor over the intricacies of the adaptation technique itself. This insight sheds light on the critical role of feature extraction in the context of few-shot semantic segmentation. This research not only contributes valuable insights into the comparative performance of vision foundation models in the realm of few-shot semantic segmentation but also highlights the significance of a robust feature extractor in this domain.
Index Terms—  Deep Learning, Foundation models, Semantic Segmentation, Few-Shot Learning
CityscapesCOCOPPD0020202020404040406060606080808080100100100100mIoUDINO V2SAMResNetCLIPMAE
Fig. 1: mIoU of various pretrained models when adapted to 3 different 1-shot segmentation tasks (only best performance achieved out of 5 tested methods is displayed).
1 Introduction
Semantic segmentation, the task of pixel-level object classification within images, i.e, assigning a class to every pixel in the image, plays a pivotal role in the landscape of computer vision applications [1]. In recent years, few-shot semantic segmentation has emerged as a challenging yet essential frontier [2]. This paradigm seeks to equip models with the ability to adapt to novel object classes with minimal labeled data. A standard pipeline for few-shot semantic segmentation typically comprises two key components: a pretrained feature extractor, often derived from a vision foundation model, and an adapter method. This setup leverages the knowledge embedded within the feature extractor and adapts it to the limited labeled data without the risk of overfitting. While extensive discussions regarding the selection of feature extractors for few-shot classification have been well-documented [3], the question of the optimal feature extractor for few-shot semantic segmentation remains largely unaddressed in the literature, to the best of our knowledge.
The primary objective of our research is to investigate which vision foundation model yields the most effective few-shot semantic segmentation pipeline. To this end, we introduce a novel benchmark, constructed using three widely-recognized semantic segmentation datasets, and systematically evaluate multiple adaptation methods. The four foundation models under consideration are DINO V2 [4], Segment Anything (SAM) [5], CLIP [6], and a Masked AutoEncoder [7] pretrained on ImageNet-1K. Our experimental results consistently demonstrate the superior performance of DINO V2 across various datasets.
In summary, our study makes the following key contributions:
  • We introduce a standardized benchmark for the evaluation of few-shot semantic segmentation in the context of vision foundation models.
  • Based on this benchmark, we conduct a comprehensive comparison of five feature extractors (four foundation models) and five adaptation methods.
Our findings indicate that DINO V2 consistently outperforms all other models across multiple datasets and adaptation methods, with SAM yielding competitive results, particularly on specific datasets. Additionally, we find that simple and efficient adaptation methods produce results comparable to more complex counterparts. We anticipate that our research will assist newcomers in selecting appropriate solutions for their few-shot semantic segmentation tasks and contribute to the rigorous benchmarking of this evolving field.
2 Benchmark
2.1 Task Formulation
In accordance with the established terminology of few-shot classification [8], our benchmark revolves around the concept of a support set $\mathcal{S}$ and a query set $\mathcal{Q}$. The support set $\mathcal{S}$ comprises a limited number of images, each accompanied by their corresponding ground truth segmentation. These support images are employed for the training or calibration of a generic semantic segmentation model. In contrast, the query set $\mathcal{Q}$ serves as the evaluation dataset on which we compute the mean Intersection-over-Union (mIoU). In practical applications, $\mathcal{Q}$ is typically not available and is utilized exclusively for benchmarking purposes.
To compute the mIoU, we calculate the Intersection-over-Union (IoU) for each class $i$ and then average these values:
$\mathrm{mIoU}=\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{TP}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}}\,,$
where $n$ represents the total number of classes or objects, $\mathrm{TP}_{i}$ is the true positive count for class $i$, $\mathrm{FP}_{i}$ is the false positive count for class $i$, and $\mathrm{FN}_{i}$ is the false negative count for class $i$. One noteworthy distinction from few-shot classification tasks is that individual images typically contain instances from various classes of interest. As such, it is challenging to define tasks in the traditional $k$-shot manner, where each class is associated with exactly $k$ elements in the support set.
In our proposed framework, we introduce a more practical approach. We define a $k$-shot segmentation task as one in which the support set $\mathcal{S}$ contains at least $k$ instances of each class. In a scenario involving a total of $n$ classes of interest, a $k$-shot task mandates the inclusion of precisely $nk$ unique images within the support set.
2.2 Benchmark Generation
To construct few-shot tasks from readily available semantic segmentation datasets, we employ the following sampling procedure:
  1. 1.
    Initialize an empty support set, denoted as $\mathcal{S}=\emptyset$.
  2. 2.
    For each class $i$ (excluding the background class) within the chosen training dataset $\mathcal{D}$, we compile a list of all training images that contain at least one instance of that class. This results in $\mathcal{D}_{i}=\{\text{image}\in\mathcal{D}\mid\text{image contains instancesof class }i\}$.
  3. 3.
    Following an arbitrary order of classes, we randomly select $k$ images from $\mathcal{D}_{i}$ without replacement to be incorporated into the support set $\mathcal{S}$. During this process, we ensure that images previously selected are excluded from consideration.
  4. 4.
    In the rare event that there are insufficient remaining images for a particular class $i$, we reset the support set to empty $\mathcal{S}=\emptyset$ and recommence the procedure from Step 1. It is important to note that this reset condition did not occur in the course of our experiments.
Fig. 2: The overall pipeline of the Linear method: we freeze the backbone and train the shallow decoder on support set, and then we test the encoder+decoder on the query set
2.3 Datasets
In constructing our proposed benchmark, we have selected three prominent semantic segmentation datasets, each offering unique characteristics and challenges:
Cityscapes [9] is a large scale dataset on semantic understanding of urban scene streets. It contains 2,975 images for training, 500 for validation and 1,525 for testing. It is annotated using 19 classes (such as “car”, “road”, “building”) and has input resolution 1024x2048.
COCO (Microsoft Common Objects in Context) [10] is a large scale dataset used for multiple tasks including semantic segmentation. The 2017 release contains 118k images for training, 5k for validation and 41k for testing. It is annotated using 80 classes (such as “person”, “bicycle”, “elephant”). Images in COCO often exhibit varying input resolutions, commonly falling between 400 and 640 pixels for both width and height.
PPD (PLANT PHENOTYPING DATASETS) [11] is a small dataset used for multiple tasks including plant segmentation (foreground to background segmentation) and leaf segmentation (multi-instance segmentation). We focus in our work on the plant segmentation task, where the dataset contains 783 images, 2 classes (foreground and background). Typical input resolutions for PPD images hover around 500 pixels for both width and height.
3 Backbones and Adaptation Methods
In this section, we introduce the methodologies that form the basis of our comparative study. Drawing from existing literature on both few-shot classification [3] and few-shot semantic segmentation [12], we have identified that the majority of methods rely on a combination of two fundamental components.
The first essential component is a pretrained model, typically a transformer or a convnet. Notably, these pretrained models are not necessarily designed specifically for semantic segmentation but offer valuable preexisting knowledge.
The second critical component is an adaptation method, a mechanism devised to tailor the knowledge encoded within the pretrained model for effective utilization in the specific task of interest.
In the upcoming subsections, we provide comprehensive details on the specific models and adaptation methods that constitute the experimental foundation of our study.
3.1 Pretrained Models
In our study, we investigate five pretrained models, each offering unique characteristics and advantages:
MAE (Masked AutoEncoder) [7] is pretrained on ImageNet-1K [13] using a self-supervised technique involving masking portions of input images and reconstructing them using partial tokens from the unmasked segments. It comprises a Vision Transformer (ViT) [14] component that maps observed image regions into latent representations and a decoder for image reconstruction. Previous research [7] has demonstrated MAE’s effectiveness in transfer learning for segmentation tasks, particularly in semantic segmentation.
SAM (Segment Anything Model) [5] is a recent foundation model for segmentation, pretrained on the SA-1B dataset, which includes 11 million images and 1 billion segmentation masks. SAM exhibits robust generalization capabilities across a broad range of segmentation tasks. It utilizes a Vision Transformer (ViT) [14] initialized with MAE [7], a prompt encoder, and a mask decoder. SAM’s image encoder generates embeddings from input images, which are subsequently employed by the mask decoder in conjunction with the prompt encoder. While SAM’s mask decoder outputs class-agnostic masks intended for segmenting instances or objects at various granularities, our specific task focuses exclusively on the encoder’s role as a feature extractor and does not require prompts.
DINO v2 [4] is a foundational vision model trained using the self-supervised DINO technique [15]. The model comprises a ViT trained on a diverse range of datasets, including notable semantic segmentation datasets like Cityscapes [9], ADE20k [16, 17], and Pascal VOC 2012 [18]. DINO v2 has previously been evaluated on semantic segmentation datasets in [4], utilizing the frozen backbone and a simple segmentation decoder.
CLIP (Contrastive Language-Image Pre-training) [6] is a model that consists of two encoders, a vision encoder, and a text encoder. These encoders are jointly trained using a contrastive loss to align embeddings of images with their corresponding captions. The model is pretrained on a private dataset of 400 million (image, text) pairs and has gained significant attention for its utility in zero-shot classification on downstream tasks. In our study, we exclusively leverage ViT version of the visual encoder component of CLIP.
FCN (Fully Convolutional Networks) [19] represents a conventional convolutional architecture extensively used for semantic segmentation tasks. It comprises an encoder and a decoder, with the former composed of convolutional layers and the latter incorporating deconvolutional layers. In our configuration, the encoder is based on ResNet50 [20] and is pretrained on a subset of COCO, specifically using the 20 classes shared with the Pascal VOC dataset. We incorporate FCN as a convolution-only benchmark for comparison.
3.2 Methods
We then consider five adaptation methods in our study:
Linear: A common approach [4] consists in freezing the pretrained backbone and adding a segmentation head, called decoder, consisting of a batch normalization layer followed by a convolutional 1x1 layer. Figure 2 illustrates the overall pipeline for this linear adaptation method.
Multilayer: It has been shown in [4] that using a concatenation of multiple layers from the pretrained model as output can boost performance in downstream tasks. We thus propose to use a linear method on the concatenation of the last 4 blocks of the pretrained model.
Singular Value Fine-tuning [21] (SVF): The SVF method decomposes parameters into three matrices using Singular Value Decomposition (SVD) and fine-tunes only the singular values while keeping the rest frozen. In [21], this approach demonstrated promising results in few-shot semantic segmentation tasks, showcasing its ability to specialize the pretrained model without overfitting. In our study, we initially apply the linear method and then proceed to fine-tune the singular values alongside the decoder parameters.
Low-Rank Adaptation [22] (LoRA): In this approach, we maintain the pretrained backbone’s frozen state and introduce low-rank trainable adapters, as described in [22]. Similar to the SVF method, LoRA enables tuning only a limited number of parameters. Our procedure begins with training the decoder using the linear method and then introduces LoRA adapters to the encoder. Subsequently, we fine-tune both the adapters and decoder parameters.
Fine-tuning: The fine-tuning method, a straightforward yet effective approach, involves fine-tuning the entire pretrained model. This method has demonstrated state-of-the-art performance in few-shot classification [3]. However, it is susceptible to overfitting due to the limited number of training samples compared to the typically extensive parameter pool in the considered encoders.
4 Experiments
4.1 Training procedure
To ensure the reproducibility of our results, we conducted all experiments with three fixed random seeds and computed the mean Intersection-over-Union (mIoU) across these runs. For preprocessing, we followed the approach outlined in [23]. This involved random horizontal flipping, random scaling of the shorter side within the range of [400, 1600] while preserving the aspect ratio, and random cropping to a size of 1024x1024. We applied a subset of RandAug [24] operations, including auto contrast, equalize, rotate, color, contrast, brightness, and sharpness. Subsequently, all images were resized to 1024x1024 resolution, except for CLIP and MAE, which accepted input sizes of 224x224. We acknowledge that the resolution variation, depending on the considered model, may introduce some analysis challenges but was an unavoidable aspect of our setup. For the learning rate, we employed the poly learning rate scheduler with a power of 0.9 for the linear and multilayer methods. Specifically, we used a learning rate of 0.2 for Cityscapes, 0.05 for COCO, and 0.001 for PPD. For fine-tuning methods (SVF, LoRA, and fine-tune) we did a grid search across models and datasets with values including $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, and $10^{-6}$ to determine the one yielding the best average performance. Linear and multilayer methods were trained with a batch size of 4. For the Fine-tuning, SVF, and LoRA methods, we initiated the training with the linear method and subsequently performed fine-tuning with the reduced batch size of 2, addressing memory limitations. Training duration extended to 200 epochs for Cityscapes and PPD, and 100 epochs for COCO. The query set consisted of the standard validation sets for Cityscapes and COCO. For PPD, given the absence of a conventional split, we performed a fixed 50/50 random split into training and validation sets. Our models were trained on a combination of NVIDIA A6000 and NVIDIA RTX3090 GPUs. Runtime varied across datasets and models, ranging from approximately 30 minutes per run for MAE, CLIP, and ResNet on Cityscapes to approximately 4 hours per run for DINO V2 on COCO. For the size of the models, we choose the base version of the vision transformers, leading to a 89M parameter model for SAM and 86M for DINO V2, CLIP and MAE. For the FCN we chose a ResNet50 as an encoder, leading to a 23M parameter model.
Table 1: mIoU averaged over our three considered datasets in 1-shot few-shot semantic segmentation.
Linear Multilayer SVF LoRA Fine-tuning
SAM 40.66$\pm$09.00 43.36$\pm$02.52 41.84$\pm$16.83 43.10$\pm$03.38 42.35$\pm$09.91
DINO V2 54.78$\pm$03.09 53.51$\pm$03.10 57.77$\pm$04.20 57.67$\pm$02.50 57.21$\pm$04.04
CLIP 35.88$\pm$10.90 39.57$\pm$07.29 40.90$\pm$13.93 38.23$\pm$04.50 38.69$\pm$10.20
MAE 35.50$\pm$10.16 36.65$\pm$07.38 37.76$\pm$08.08 36.37$\pm$04.22 36.46$\pm$10.28
FCN-ResNet50 37.78$\pm$04.17 44.08$\pm$02.16 43.40$\pm$08.63 41.45$\pm$03.70 39.63$\pm$06.49
4.2 Main results
In this section, we aim to identify the optimal combination of pretrained models and adaptation methods based on mIoU. Table 1 presents the average performance on the three considered datasets for all possible combinations of pretrained models and adaptation methods, along with their standard deviation, in the context of 1-shot problems.
From this initial experiment, several conclusions can be drawn:
  • DINO V2 consistently outperforms all other models across various settings. This superiority is especially notable on Cityscapes and COCO, where DINO V2 surpasses other models by a significant margin (see Appendix Table 2 for the detailed results). It is worth noting that Cityscapes is included in the training images of DINO V2, although it only had access to the images and not their associated segmentation masks.
  • SAM encoder exhibits strong performance on average. It outperforms CLIP, MAE, and ResNet on both Cityscapes and PPD. However, it yields comparatively poor results on COCO (refer to Appendix Table 2). This difference is likely attributed to our decision to retain the SAM image encoder while discarding the original decoder, which was jointly trained with the image and prompt encoders.
  • The multilayer method outperforms the linear approach for most models, indicating that including earlier feature maps, which are more locally focused [25, 14], provides the decoder with greater segmentation flexibility. Furthermore, it exhibits a lower standard deviation, indicating greater stability across different runs.
  • Despite variations in the number of training parameters (less than 1% for LoRA, around 2% for SVF, and 100% for finetune), the finetuning methods (SVF, LoRA, and Finetune) yield comparable results.
5 Related Work
In this section, we discuss related work in the literature in few-shot classification and semantic segmentation.
5.1 Few-Shot Classification
Few-shot classification initially focused on in-domain tasks, with benchmarks involving dataset splitting into training, validation, and fake few-shot task generation for evaluation [26, 8, 27, 28]. The introduction of Meta-dataset [29] marked a shift towards the cross-domain setting, where the dataset used for model training differs from the tasks evaluated later, aligning with more realistic scenarios. More recently, the field has embraced the use of foundation models, adapting them to few-shot tasks. Notably, benchmarks such as those in [30] explore the application of CLIP [6] to 11 downstream tasks.
Methodologically, diverse approaches have emerged, ranging from hallucination-based [31, 32] to meta-learning [28, 33, 34, 35]. Straightforward methods have also demonstrated effectiveness [36, 37, 38]. Recent surveys, such as [3], suggest that optimal approaches often involve simple finetuning atop competitive pretrained models, augmented with an additional linear layer. This observation motivates the methods introduced in our paper.
5.2 Semantic Segmentation and Few-Shot Semantic Segmentation
Few-shot semantic segmentation is not a recent topic. Earlier works, such as [2] and [39], explored multiple benchmarks, including PASCAL-$5^{i}$ and COCO-$20^{i}$, reminiscent of in-domain benchmarks used in few-shot classification. Subsequent research ventured into using pretrained models on ImageNet [40], resembling the Meta-dataset approach in few-shot classification. Notably, most of these studies assumed binary labelling for the few-shots, emphasizing a single object of interest.
In contrast, our approach considers that the provided few-shots are fully labelled, and we leverage pretrained foundation models, not limited to those trained solely on ImageNet. This approach aligns more closely with contemporary industrial requirements, focusing on the utilization of the most efficient available models.
6 Conclusion
In conclusion, DINO V2 emerges as the top-performing backbone model, consistently outperforming other foundation models. While SAM exhibits strong overall results, it presents unexpected limitations on certain datasets. Surprisingly, our experiments underscore the competitive advantages of straightforward methods such as simple linear heads trained on the few given shots, challenging the necessity of complex procedures.
This work not only presents a comparison of techniques but also a practical benchmark. We hope this benchmark will contribute to improving few-shot semantic segmentation methods and will provide a solid ground for comparison purposes.